OpenStudy (ajanijones):

The table shows the results of drawing one marble from a bag of colored marbles. Each marble was replaced after the outcome was recorded. If there are a total of 50 marbles in the bag, about how many of the marbles are black? Color Number of Pulls red 3 green 7 orange 4 black 6 A. 6 B. 12 C. 9 D. 15

OpenStudy (anonymous):

6 / (3+7+4+6) * 50 = 6/20 * 50 = 15 black